OpenStudy (anonymous):

Use a sum or difference formula to find the exact value of the given expression. cos 5π/12

OpenStudy (freckles):

\[\frac{5\pi}{12}=\frac{\pi}{6}+\frac{\pi}{4}\] or if it easier convert to degs then find the special sum \[\frac{5\pi}{12} rad=\frac{5 \pi}{12} \cdot \frac{180 \deg}{\pi rad}= \frac{5(180)}{12} \deg=5(15) \deg=75 \deg \\ 75 \deg=30\deg+45 \deg \]

OpenStudy (freckles):

now you just have to use the cos sum identity

OpenStudy (anonymous):

THANKS